# QSpec: Speculative Decoding with Complementary Quantization Schemes

### v1

@@ 1. Q1:
    Does QSpec serve as a general solution without compromising quality compared to high-precision weight-only quantization methods? (Sec. 4.2)
@@ 2. Q2:
    Does QSpec consistently accelerate high-precision weight-only quantization methods? (Sec. 4.3)
@@ 3. Q3:
    What is the common acceptance rate of QSpec, and how does the length of draft tokens affect it? (Sec. 5)
@@ 4.1 General Setup
Benchmarks. We assess QSpec with two primary criteria: (1) generation fidelity and (2) end-to-end serving speedup. For fidelity evaluation, we adopt not only traditional tasks, including PIQA (500, 10-shot) (Bisk et al., 2020), WinoGrande (500, 5-shot) (Sakaguchi et al., 2019), and WikiText2 (Merity et al., 2016), but also challenging multi-step reasoning tasks such as GSM8K (All, 8-shot) (Cobbe et al., 2021), MATH (All, 4-shot) (Hendrycks et al., 2021), MBPP (200, 0-shot) (Austin et al., 2021), and HumanEval (All, 0-shot) (Chen et al., 2021). To measure the acceleration, we use all the above reasoning tasks and two additional chatbot datasets, namely ShareGPT (RyokoAI, 2021) and LMsys-1K (Zheng et al., 2023). Following the setup of Atom (Zhao et al., 2024c), we randomly sampled the dataset for the request prompts to reduce the workload. Due to memory limitations, we vary the batch size from 8 to 32 and serve all requests in a first-come, first-served (FCFS) manner. Once any request is finished, we refill the batch, adhering to the continuous batching approach of ORCA (Yu et al., 2022). We use greedy sampling for token generation.
Base Models. To assess the effectiveness and scalability of our approach, we conduct experiments using multiple models from the Llama family (Dubey et al., 2024) [^1] https://www.huggingface.co/meta-llama with varying scales and capacities: Llama3.2-3b, Llama2-7b, Llama3-8b-instruct, and Llama2-13b.
Implementation.
All experiments are performed on a node equipped with four NVIDIA A100 GPUs (40GB HBM each) running CUDA 12.5. To demonstrate the versatility of QSpec, we implement two SOTA 4-bit quantization methods, namely Atom (Zhao et al., 2024c) and QuaRot (Ashkboos et al., 2024). For W4A16 configurations, we incorporate AWQ-style (Lin et al., 2024a) weight dequantization logic for runtime inference. We select Atom to showcase the acceleration of QSpec. We use these Group-wise quantization schemes with a group size of 128. With the draft token length $\gamma$ as 3, we simulate the performance of QSpec by initially employing fake quantization to fully emulate the execution flow, encompassing both the draft and verify stages of QSpec. Subsequently, we replay the collected traces with real kernel execution to accurately reproduce the latency. [^2] Atom’s kernel only supports shape-specific models. We modify the model structure to meet requirements while maintaining the original model size.
@@ 4.2 Fidelity Evaluation
Table 3: Performance of different quantization methods across multiple general and reasoning benchmarks: PIQA, WinoGrande, GSM8K, MATH, MBPP, and HumanEval. The quality degradation ratio is calculated by $1-\frac{W4A4}{W4A16}$.
Method Quantization WikiText-255footnotemark: 5 PIQA WinoGrande GSM8K MATH MBPP HumanEval
@@
Atom W4A16 7.87 74.8 62.0 73.4 24.3 41.5 52.4
QSpec 7.87 75.0 62.0 73.4 24.3 39.0 52.4
W4A4 8.6 (-9.58%) 65.8 (-12.03%) 56.2 (-9.35%) 54.7 (-25.47%) 15.5 (-36.21%) 30.5 (-26.50%) 31.7 (-39.50%)
QuaRot W4A16 8.58 74.2 59.4 70.5 25.1 40.0 45.7
@@
W4A4 10.2 (-19.24%) 62.6 (-15.63%) 53.8 (-9.43%) 42.0 (-40.43%) 12.3 (-51.11%) 28.5 (-28.75%) 28.0 (-38.73%)
QSpec effectively maintains the generation quality of W4A16, whereas W4A4 does not. As listed in Table 3, with the draft verification of W4A16, QSpec exhibits only minimal performance fluctuations compared to W4A16. This negligible variation may stem from the nondeterministic algorithms of PyTorch [^3] https://pytorch.org/docs/stable/notes/randomness.html or occasional cases where two tokens have the same maximum prediction probability. In contrast, W4A4 experiences a substantial performance decline exceeding 10% across most tasks, with the reduction becoming more pronounced as task difficulty increases. For instance, compared to GSM8K and MBPP, the performance drop for W4A4 is much greater on the more challenging MATH and HumanEval tasks, showing declines of 51.11% and 38.73%, respectively. On the other hand, this also highlights the higher sensitivity of multi-step reasoning tasks to the negative effects of quantization compared to regular tasks, such as WikiText-2444To measure perplexity, only a single prefill stage is necessary, bypassing the verify stage. and WinoGrande. This observation aligns with our earlier analysis in Sec. 2, encouraging incorporating multi-step reasoning tasks into quantization evaluation.
@@ 4.3 Acceleration Evaluation
Table 4: Comparison of token generation throughput across different model sizes, quantization configurations, and batch sizes for various datasets. All values are measured in token/s. “Avg.” denotes the average speedup ratio for the corresponding row or column.

### v2

@@ 1. Q1:
    Does QSpec preserve the quality of high-precision weight-only quantization? (Sec. 4.2)
@@ 2. Q2:
    Does QSpec accelerate high-precision weight-only quantization methods and surpass SOTA speculative decoding method in quantized scenarios? (Sec.4.3)
@@ 3. Q3:
    How do various factors (e.g., quantization method, sequence length) influence QSpec’s acceptance rate and acceleration performance? (Sec. 4.4)
@@ 4.1   General Setup
Benchmarks. We assess QSpec with two primary criteria: (1) generation fidelity and (2) end-to-end serving speedup. For fidelity evaluation, we adopt not only traditional tasks, including PIQA (500, 10-shot) (Bisk et al., 2020), WinoGrande (500, 5-shot) (Sakaguchi et al., 2019), and WikiText2 (Merity et al., 2016), but also challenging multi-step reasoning tasks such as GSM8K (All, 8-shot) (Cobbe et al., 2021), MATH (All, 4-shot) (Hendrycks et al., 2021), MBPP (200, 0-shot) (Austin et al., 2021), and HumanEval (All, 0-shot) (Chen et al., 2021). To measure the acceleration, we use all the above reasoning tasks and two additional chatbot datasets, namely ShareGPT (RyokoAI, 2021) and LMsys-1K (Zheng et al., 2023). Due to space constraints, we present results for GSM8K, HumanEval, and LMsys-1K in the main text, with remaining results detailed in the Appendix A. Following the setup of Atom (Zhao et al., 2024b), we randomly sampled the dataset for the request prompts to reduce the workload. Due to memory limitations, we vary the batch size from 8 to 32 and serve all requests in a first-come, first-served (FCFS) manner. Once any request is finished, we refill the batch, adhering to the continuous batching approach of ORCA (Yu et al., 2022). All experiments use greedy sampling for token generation.
Models. To assess the effectiveness and scalability of our approach, we conduct experiments using multiple models from the Llama family (Touvron et al., 2023; Dubey et al., 2024) with varying scales and capacities: Llama3.2-3b, Llama2-7b, Llama3-8b-instruct, and Llama2-13b.
Implementations. All experiments are performed on a node equipped with four NVIDIA L20 GPUs (48GB HBM each) running CUDA 12.5. To demonstrate the versatility of QSpec, we implement two SOTA 4-bit quantization methods, namely Atom (Zhao et al., 2024b) and QuaRot (Ashkboos et al., 2024). For W4A16 configurations, we incorporate AWQ-style (Lin et al., 2024a) weight dequantization logic for runtime inference. We select Atom to showcase the acceleration of QSpec. We use these Group-wise quantization schemes with a group size of 128. With the draft token length $\gamma$ as 3, we simulate the performance of QSpec by initially employing fake quantization to fully emulate the execution flow, encompassing both the draft and verify stages of QSpec. Subsequently, we replay the collected traces with real kernel execution to accurately reproduce the latency.
Baselines. We evaluate QSpec against baseline quantization configurations: W4A4, W4A16, and W16A16. For comparison with SOTA speculative decoding approaches, we include EAGLE (Li et al., 2024a), which employs a pruned draft model with tree-structured speculative decoding. In our quantized experiments, we utilize FP16 precision for the EAGLE draft model and EAGLE-Quant (W4A16) for the target model. This choice was necessitated by two factors: (1) the official EAGLE quantization implementation (fast-gpt) lacks efficient batching support, and (2) applying GPTQ quantization to the EAGLE draft model resulted in substantial degradation of the acceptance rate.
@@ 4.2   Fidelity Evaluation
Table 3: Performance of different quantization methods across multiple general and reasoning benchmarks: PIQA, WinoGrande, GSM8K, MATH, MBPP, and HumanEval. The quality degradation ratio is calculated by $\frac{W4A4}{W4A16}-1$.
Method Quantization WikiText-2 PIQA WinoGrande GSM8K MATH MBPP HumanEval
@@
Atom W16A16 7.73 76.8 61.4 76.2 24.9 42.5 53.0
W4A16 7.87 74.8 62.0 73.4 24.3 42.0 52.4
QSpec 7.87 75.0 62.0 73.4 24.3 40.5 52.4
W4A4 8.6 (+9.58%) 65.8 (-12.03%) 56.2 (-9.35%) 54.7 (-25.47%) 15.5 (-36.21%) 33.0 (-21.43%) 31.7 (-39.50%)
QuaRot W16A16 7.73 76.8 61.4 76.2 24.9 42.5 53.0
W4A16 8.58 74.2 59.4 70.5 24.7 40.0 45.7
@@
W4A4 10.2 (+19.24%) 62.6 (-15.63%) 53.8 (-9.43%) 42.0 (-40.43%) 12.3 (-51.11%) 28.5 (-28.75%) 28.0 (-38.73%)
QSpec effectively maintains the generation quality of W4A16, whereas W4A4 does not. As listed in Table 3, with the draft verification of W4A16, QSpec exhibits only minimal performance fluctuations compared to W4A16. This negligible variation may stem from the nondeterministic algorithms of PyTorch (PyTorch Contributors, 2024) or occasional cases where two tokens have the same maximum prediction probability. In contrast, W4A4 experiences a substantial performance decline exceeding 10% across most tasks, with the reduction becoming more pronounced as task difficulty increases. For instance, compared to GSM8K and MBPP, the performance drop for W4A4 is much greater on the more challenging MATH and HumanEval tasks, showing declines of 51.11% and 38.73%, respectively. On the other hand, this also highlights the higher sensitivity of multi-step reasoning tasks to the negative effects of quantization compared to regular tasks, such as WikiText-2and WinoGrande. This observation aligns with our earlier analysis in Sec. 2, encouraging incorporating multi-step reasoning tasks into quantization evaluation.
@@ 4.3   Acceleration Evaluation
Table 4: Comparison of token generation throughput across different model sizes, quantization configurations, and batch sizes for various datasets. All values are measured in token/s. ”Avg.” denotes the average speedup ratio for the corresponding row or column. $\dagger$” indicates the failure of W4A16 kernels to support the configuration. W16A16 results (when OOM) are collected using pipeline parallelism Harlap et al. (2018) with multiple cards.
Model Method Batch GSM8K HumanEval LMsys-1k
3B W16A16 8 511.1 647.2 711.2
16 666.5 948.3 1126.4
32 833.4 $1$.6 1553.3
W4A4 8 804.7 892.6 990.3
16 1109.1 1289.8 1581.0
32 1424.3 1488.2 2194.4
W4A16 8 420.0 535.7 559.8
16 578.5 804.4 925.8
32 726.3 954.4 1336.4
QSpec 8 594.1 (1.41$\times$) 723.6 (1.35$\times$) 738.8 (1.32$\times$)
16 811.5 (1.40$\times$) 1042.1 (1.30$\times$) 1171.4 (1.27$\times$)
32 1030.4 (1.42$\times$) 1248.5 (1.31$\times$) 1576.0 (1.18$\times$)
Avg. 1.41$\times$ 1.32$\times$ 1.25$\times$
7B W16A16 8 213.4 316.7 285.3
16 290.3 505.1 441.6
32 340.9 663.6 564.2
W4A4 8 349.5 471.2 419.4
16 496.6 749.5 642.6
32 620.0 1043.9 865.5
W4A16 8 165.0 240.2 220.2
16 231.8 407.3 358.0
32 268.9 555.9 470.1
QSpec 8 253.7 (1.54$\times$) 350.9 (1.46$\times$) 310.3 (1.41$\times$)
16 359.8 (1.55$\times$) 555.2 (1.36$\times$) 473.1 (1.32$\times$)
32 441.8 (1.64$\times$) 749.4 (1.35$\times$) 628.4 (1.34$\times$)
Avg. 1.58$\times$ 1.39$\times$ 1.36$\times$
13B W16A16 8 121.9 182.0 160.1
16 169.6 291.0 243.0
32 202.4 423.5 334.2
W4A4 8 194.7 261.5 228.2
16 288.3 424.9 348.4
32 369.8 665.4 508.8
W4A16 8 94.8 140.0 127.9
16 136.1 236.9 207.2
32 $\dagger$ 365.5 287.4
QSpec 8 148.2 (1.56$\times$) 201.2 (1.44$\times$) 174.0 (1.36$\times$)
16 212.8 (1.56$\times$) 323.3 (1.36$\times$) 266.9 (1.29$\times$)
32 266.6 ($\dagger$) 483.0 (1.32$\times$) 379.3 (1.32$\times$)
Avg. 1.56$\times$ 1.37$\times$ 1.32$\times$
QSpec exhibits a substantial efficiency boost compared to W4A16. In Table 4, we present the token generation throughput for both QSpec and W4A16 across different model sizes, quantization configurations, and batch sizes on diverse datasets. On average, QSpec achieves a throughput increase of 1.38×1.38$\times$1.38 × over W4A16 across all settings, with a peak improvement of 1.64×1.64$\times$1.64 ×.
Larger models tend to yield better speedup ratios. We find a gradually-increasing acceleration as the base model scales up. This overall upward trend indicates a promising outlook for our approach with larger models, although further experiments are needed for confirmation. Due to resource limitations, this will be addressed in future work.
Figure 5: Per-valid-token latency decomposition for different methods. The latency of QSpec is further decomposed into draft and verify categories for details.
Latency composition. As illustrated in Figure 5, we compute the per-valid-token latency by dividing the total latency by only the number of accepted tokens before averaging on all evaluation datasets. Notably, QSpec achieves remarkable latency savings ranging from 28.5% to 39.7%. Besides, the per-token latency is further decomposed into two parts: draft and verify. Clearly, the primary gains of QSpec arise from the rapid drafting capability and the reduced latency achieved through the parallel verification of multiple tokens.
Table 5: Performance comparison of EAGLE-Quant, QSpec, W4A16, and W4A4 on Llama-2-7b-chat-hf across different batch sizes and benchmarks. “OOM” indicates out-of-memory. Better cases for QSpec or EAGLE is marked in gray. In the case of batch size=8, the speedup ratio of QSpec compared to EAGLE is indicated in parentheses next to the data points.
Method Batch Size GSM8K (8-shot) HumanEval (0-shot) LMsys-1k
EAGLE 1 65.81 49.15 71.29
8 140.16 136.86 167.57
16 OOM OOM OOM
1 51.25 54.22 56.14
QSpec 8 208.95 (1.49$\times$) 185.99 (1.36$\times$) 260.48 (1.55$\times$)
16 292.82 255.11 463.35
W4A16 1 59.80 72.04 72.27
8 146.34 163.54 213.66
16 190.09 211.49 371.38
W4A4 1 64.79 73.47 71.55
8 284.84 256.54 393.12
16 401.77 330.71 713.67
Memory and batching efficiency over speculative decoding. Table 5 compares QSpec and EAGLE on Llama-2-7b-chat-hf. EAGLE delivers optimal performance for single-sequence inputs (batch size=1). However, its efficiency degrades as the batch size grows (8 and 16). This observation aligns with our earlier analysis in Sec. 3.2: EAGLE’s tree-structured drafting mechanism, designed to reconcile discrepancies between the draft and target models, introduces additional latency that reduces the gains from higher acceptance rates in batched serving under quantization. Furthermore, the increasing key-value (KV) storage of EAGLE’s draft model leads to out-of-memory (OOM) issues at batch size 16. In contrast, QSPEC demonstrates superior scalability and memory efficiency.
4.4   Ablation Study
Ablation on draft token length. To assess parameter sensitivity, we vary the draft token lengths $\gamma$, the sole hyper-parameter of QSpec, from 2 to 6 across all the benchmarks using Llama3.2-3b and Llama3-8b-instruct models. As depicted in Figure 6, an increase in $\gamma$ leads to a gradual decline in the token acceptance rate, since all subsequent tokens are discarded once a token is rejected. Nevertheless, even at $\gamma=6$, the token acceptance rate remains relatively high, approximately $74\%$, compared to $28\sim 58\%$ in 160m-7b draft-target model pair under $\gamma=5$ in conventional speculative decoding (Liu et al., 2024). Additionally, a consistent improvement in throughput is observed compared to W4A16, indicating the robustness of QSpec with respect to $\gamma$. More comprehensive comparison is provided in Appendix A.2.
Table 6: Ablation study comparing acceptance rates (%) across base quantization methods using QSpec.
Quantization Method ShareGPT MATH (4-shot) MBPP (0-shot)
Atom 83.8 89.4 88.6
QuaRot 81.6 88.9 85.4
Ablation on base quantization method. As shown in Table 6, QSpec consistently achieves high acceptance rates across diverse quantization methods and datasets.
5   Related work
Quantization is a common technique for deploying LLMs on resource-limited scenarios. Broadly, recent quantization algorithms can be classified into two categories: weight-only W4A16 and weight-activation joint W4A4. Notably, AWQ (W4A16) (Lin et al., 2024a) redistributes the quantization burden by scaling salient weight channels to protect them from degradation. In contrast, W4A4 aggressively quantizes activations to leverage low-precision hardware for improved speed at the cost of model quality degradation. To address this challenge, Atom (Zhao et al., 2024b) proposes reordering outlier channels in the activation through offline profiling. Similarly, QuaRot (Ashkboos et al., 2024) employs Hadamard matrices to apply computational invariance on weights. Despite these advancements, our observations indicate that W4A4 methods still exhibit substantial degradation compared to weight-only quantization approaches across multi-step reasoning tasks. Works such as W4A8 (Lin et al., 2024b) and adaptive quantization (Zhao et al., 2024a; Dong et al., 2024) seek to identify an optimal trade-off point. However, these methods struggle to fully preserve the generation quality associated with higher precision.
Speculative Decoding leverages a draft model to generate candidate tokens, which are then validated by a target model (Leviathan et al., 2023). Recent research has primarily focused on improving the acceptance rate and generation speed of candidate tokens. SpecInfer (Miao et al., 2024) introduces a boost-tuned small language model to generate candidate tokens in tree structures, enabling single-pass verification. In contrast, EAGLE (Li et al., 2024a) adopts an aggressive pruning strategy for the draft model’s architecture, allowing penultimate layer feature prediction with minimal computational overhead. Self-speculative decoding, a subset of this technique, employs a single model for both draft generation and verification. LayerSkip (Elhoushi et al., 2024) introduces a training methodology for early exit with layer drop, subsequently verifying partially generated tokens through full model inference. Medusa (Cai et al., 2024) augments the original LLM with additional heads atop the final hidden state while relaxing the acceptance policy. However, these approaches inevitably require retraining of the original model, which can be computationally expensive and time-consuming. We further demonstrate their deficiency in batched serving under quantization scenario.
6   Conclusion
In this paper, we begin by validating that multi-step reasoning tasks can capture performance degradation incurred by activation quantization more sensitively and consistently than current evaluation protocols, advocating for their incorporation for a more comprehensive assessment. With observed high token-level similarities, we introduce QSpec, a novel quantization paradigm that seamlessly synergizes two complementary weight-shared quantization schemes with speculative decoding. Empirically, QSpec achieves up remarkable acceleration without any quality sacrifice across diverse settings and baselines. Alongside consistent memory consumption and a plug-and-play property, these advantages distinguish QSpec from any existing solution, promising it for high-fidelity quantization deployment, particularly in memory-constrained scenarios.
7   Impact Statement
This paper introduces QSpec, a novel quantization paradigm that synergistically combines complementary quantization schemes for speculative decoding to enhance computational efficiency while preserving model fidelity. The impact of QSpec is twofold.
From an academic perspective, QSpec establishes a new paradigm that decouples speed acceleration from quality preservation—a longstanding trade-off in prior quantization research. By enabling independent optimization of these objectives, our work addresses critical challenges in quantization design—particularly under modern training regimes where improved model convergence and higher parameter utilization efficiency amplify the severity of precision degradation caused by quantization  Kumar et al. (2024).
For industry applications, QSpec provides a practical pathway to accelerate inference without compromising output quality through efficient low-precision kernels. This advancement challenges hardware vendors to reconsider architectural support for low-precision execution, including specialized instruction set architectures (ISAs) and memory subsystems. The plug-and-play nature of QSpec further facilitates seamless integration into existing deployments of quantized models on edge devices, lowering barriers to adoption.
We identify no direct adverse societal impacts, as our work focuses on optimizing computational efficiency for language model serving systems. By reducing resource requirements, QSpec could broaden access to high-performance AI infrastructure, thereby contributing to more sustainable and scalable deployments of language technologies.
Appendix A Full Experiment Result.
A.1   Acceleration Evaluation
Versus quantization configurations.
Table 7: Comparison of token generation throughput across different model sizes, quantization configurations, and batch sizes for various datasets. All values are measured in token/s. “Avg.” denotes the average speedup ratio for the corresponding row or column. $\dagger$” indicates the failure of W4A16 kernels to support these batch sizes together with long sequences and the large models.